# Geodesic loops and orthogonal geodesic chords without self-intersections

### v1

@@ 3. Geodesics with product boundary conditions
Therefore this argument shows that for points $p,q$ – where here we can also allow $p=q$ – the set $\mathcal{G}^{\prime}_{\{p\}\times\{q\}}[L]$ of Riemannian metrics for which all geodesic segments of length $\leq L$ joining $p$ and $q$ are non-degenerate and prime is an open and dense subset. If now $x\in DIP_{1}(L),x\not\in\{p,q\}$ we can choose $\eta\in(0,\mathrm{inj}/3)$ such that the following holds: If a geodesic $c_{j}$ for some $j=1,\ldots,k$ enters the geodesic ball $B_{p}(2\eta)$ around $p\in DIP(L)$ with radius $2\eta,$ i.e. $c_{j}(s)\in B_{p}(2\eta),$ then there is $s_{1}$ with $c_{j}(s_{1})=p$ and $d(c_{j}(s),c_{j}(s_{1}))=|s-s_{1}|\|c_{j}^{\prime}\|\leq 2\eta.$ Hence one can use the perturbation result [18, Lem.3] resp. Lemma 2 and perturb the metric in the ball $D_{p}(2\eta)$ for all finitely many points in $DIP_{1}(L)$ such that we obtain a metric $g_{1}\in\mathcal{G}_{\{p\}\times\{q\}}^{*}(M)[L]$ in any arbitrary small neighborhood of the original metric. ∎
@@ 4. Proofs of the statements in section 1
Since $M$ is simply-connected and compact there is a number $l\in\{2,\ldots,n-1\},$ such that the sequence $b_{2lk}=\dim H_{2lk}\left(\Omega_{\{p\}\times\{q\}}(M);\mathbb{Q}\right),k\geq 1$ satisfies $b_{2lk}\geq 1.$ It was shown by Gromov th,at there is a constant $C(g)>0$ such that the critical value $\mathrm{cr}(w_{k})$ of a non-trivial homology class $w_{k}\in H_{2lk}\left(\Omega_{\{p\}\times\{q\}}(M);\mathbb{Q}\right)$ satisfies $\mathrm{cr}(w_{k})\leq C(g)k$ for all $k\geq 1,$ cf. Gromov [9], [17, Thm.5.1]. By Theorem 1 we obtain for a generic metric a sequence $c_{k}$ of simple and non-degnerate geodesic segments joining $p$ and $q$ of index $2lk$ and length $\leq C(g)k.$ This follows from the Morse inequalities since the energy-functional $E:\Omega_{\{p\}\times\{q\}}(M)\longrightarrow\mathbb{R}$ is a Morse-function. ∎
@@
As remarked in [7, p.117] the absence of orthogonal-tangent geodesic chords is an open condition with respect to the $C^{1}$-topology, hence $\mathcal{G}_{1}$ is an open subset. For a rotationally symmetric metric on the disc there are no orthogonal-tangent geodesic chords since any geodesic segment starting orthogonally from $\partial D^{n}$ will end orthogonally at the boundary $\partial D^{n}.$ Hence $\mathcal{G}_{1}$ is non-empty. Let $g_{0}$ be the Euclidean inner product on the $n$-disc $D^{n}\subset\mathbb{R}^{n},$ then let $g\in\mathcal{G}(D^{n})\longmapsto D(g)\in R^{+}$ be the smallest positive number such that $D(g)^{-1}g_{0}(v,v)\leq g(v,v)\leq D(g)g_{0}(v,v)$ for any tangent vector $v\in T_{x}D^{n},x\in D^{n}.$ Since the mapping $D:\mathcal{G}(D^{n})\longrightarrow\mathbb{R}^{+}$ is continuous the set $\mathcal{G}_{0}$ of Riemannian metrics for which all orthogonal geodesic chords of length $\leq\pi D(g)$ are non-degenerate and simple is an open and dense subset by Theorem 4(b), compare the argument in [19, Sec.6]. Then the intersection $\mathcal{G}_{2}=\mathcal{G}_{0}\cap\mathcal{G}_{1}$ contains all metrics $g$ which do not have an orthogonal-tangent geodesic chord and for which all orthogonal geodesic chords of length $\leq D(g)$ are simple and non-degenerate and do not intersect each other. Then we use the result by Giambò, Giannoni and Piccione [7, Thm.1.1] which ensures the existence of $n$ geometrically distinct orthogonal geodesic chords for a metric without orthogonal-tangent geodesic chords. ∎

### v2

@@ 3. Geodesics with product boundary conditions
Therefore this argument shows that for points $p,q$ (here we also allow $p=q$) the set $\mathcal{G}^{\prime}_{\{p\}\times\{q\}}[L]$ of Riemannian metrics for which all geodesic segments of length $\leq L$ joining $p$ and $q$ are non-degenerate and prime is an open and dense subset. If now $x\in DIP_{1}(L),x\not\in\{p,q\}$ we can choose $\eta\in(0,\mathrm{inj}/3)$ such that the following holds: If a geodesic $c_{j}$ for some $j=1,\ldots,k$ enters the geodesic ball $B_{p}(2\eta)$ around $p\in DIP(L)$ with radius $2\eta,$ i.e. $c_{j}(s)\in B_{p}(2\eta),$ then there is $s_{1}$ with $c_{j}(s_{1})=p$ and $d(c_{j}(s),c_{j}(s_{1}))=|s-s_{1}|\|c_{j}^{\prime}\|\leq 2\eta.$ Hence one can use the perturbation result [23, Lem.3] resp. Lemma 2 and perturb the metric in the ball $D_{p}(2\eta)$ for all finitely many points in $DIP_{1}(L)$ such that we obtain a metric $g_{1}\in\mathcal{G}_{\{p\}\times\{q\}}^{*}(M)[L]$ in any arbitrary small neighborhood of the original metric. ∎
@@ 4. Proofs of the statements in section 1
Since $M$ is simply-connected and compact the following holds: There is a number $l\in\{2,\ldots,n-1\},$ such that the sequence $b_{2lk}=\dim H_{2lk}\left(\Omega_{\{p\}\times\{q\}}(M);\mathbb{Q}\right),k\geq 1$ satisfies $b_{2lk}\geq 1.$ It was shown by Gromov [11] and Paternain [22, Thm.5.1] that there is a constant $C(g)>0$ such that the critical value $\mathrm{cr}(w_{k})$ of a non-trivial homology class $w_{k}\in H_{2lk}\left(\Omega_{\{p\}\times\{q\}}(M);\mathbb{Q}\right)$ satisfies $\mathrm{cr}(w_{k})\leq C(g)k$ for all $k\geq 1.$ By Theorem 1 we obtain for a generic metric $g$ a sequence $c_{k}$ of simple and non-degenerate geodesic segments joining $p$ and $q$ of index $2lk$ and length $\leq C(g)k.$ This follows from the Morse inequalities since the energy-functional $E:\Omega_{\{p\}\times\{q\}}(M)\longrightarrow\mathbb{R}$ is a Morse-function. ∎
Note that a geodesic loop $c:[0,1]/\{0,1\}\longrightarrow M$ with $p=c(0)$ is prime, if there is no $t\in(0,1)$ with $p=c(t).$ Going through the Proof of Theorem 4 and of Corollary 1 we find that in case of a surface, i.e. for $n=2,$ we obtain the following result. Here we essentially use the Perturbation Lemma 3:
Proposition 1.
Let $M$ be a compact surface, i.e. $n=\dim M=2$ and let $p\in M$ be a point.
  • (a)
    For a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all geodesic loops based at $p$ are non-degenerate and prime.
  • (b)
    For a $C^{k}$-generic Riemannian metric on the $2$-sphere $S^{2}$ the number $N_{g}(p;t)$ of geodesic loops based at $p$ satisfies $\liminf_{t\to\infty}(N_{g}(p;t)/t)>0.$
@@
As remarked in [9, p.117] the absence of orthogonal-tangent geodesic chords is an open condition with respect to the $C^{1}$-topology, hence $\mathcal{G}_{1}$ is an open subset. For a rotationally symmetric metric on the disc there are no orthogonal-tangent geodesic chords since any geodesic segment starting orthogonally from $\partial D^{n}$ will end orthogonally at the boundary $\partial D^{n}.$ Hence $\mathcal{G}_{1}$ is non-empty. Let $g_{0}$ be the Euclidean inner product on the $n$-disc $D^{n}\subset\mathbb{R}^{n},$ then let $g\in\mathcal{G}(D^{n})\longmapsto D(g)\in R^{+}$ be the smallest positive number such that $D(g)^{-1}g_{0}(v,v)\leq g(v,v)\leq D(g)g_{0}(v,v)$ for any tangent vector $v\in T_{x}D^{n},x\in D^{n}.$ Since the mapping $D:\mathcal{G}(D^{n})\longrightarrow\mathbb{R}^{+}$ is continuous the set $\mathcal{G}_{0}$ of Riemannian metrics for which all orthogonal geodesic chords of length $\leq\pi D(g)$ are non-degenerate and simple is an open and dense subset by Theorem 4(b), compare the argument in [24, Sec.6]. Then the intersection $\mathcal{G}_{2}=\mathcal{G}_{0}\cap\mathcal{G}_{1}$ contains all metrics $g$ which do not have an orthogonal-tangent geodesic chord and for which all orthogonal geodesic chords of length $\leq D(g)$ are simple and non-degenerate and do not intersect each other. Then we use the result by Giambò, Giannoni and Piccione [9, Thm.1.1] which ensures the existence of $n$ geometrically distinct orthogonal geodesic chords for a metric without orthogonal-tangent geodesic chords. ∎